Assume that the Habib Bank of Pakistan has total reserves of Rs. 50 Million. Assume also that required reserves are 15 percent of checking deposits and that bank hold no excess reserves and households hold no currency.

a. Show a T-account for Habib Bank (1 Mark)
b. Calculate the money multiplier and Calculate the money supply (1 Mark)
If the State Bank of Pakistan now raises required reserves to 25 percent of deposits,
c. What will be the effect on money multiplier? (1 Mark)
d. What will be the effect on Reserves? What will be the effect on money Supply? (1
1
Expert's answer
2021-09-09T11:34:43-0400

a.

T-account

Dr side: Assets

Required reserves 15% of 50 million 7.5 million

reserves bal. 42.5 million

balancing figure 50 million

Credit side: Liabilities

bal. carried down 50 million

b.

money multiplier = "\frac{1}{reserve.ratio}"

reserve ratio = 15% which is equal to 0.15

Therefore money multiplier = "\frac{1}{0.15}"

money multiplier = 6.7


Money Supply = Change in Reserves "\times" Money Multiplier

money supply = 50,000,000 "\times" 6.7

money supply = 335 million

c.

New reserve ratio = 0.25

New money multiplier = "\frac{1}{0.25}"

money multiplier = 4

Effect: the money multiplier will decrease to 4 from 6.7

d.

Effect: Reserves will increase by 10% of the total deposits

new money supply = Change in Reserves "\times" Money Multiplier

new money supply = 50,000,000 "\times" 4

new money supply 200 million

Effect: money supply will reduce to 200 million from 335 million